a) 200.0 g of Ca(C2H3O2)2 was found in the chemistry laboratory. 

      Calculate

          i. the molecular mass of 1 mole Ca(C2H3O2) 2.

         ii.  the number of  moles of Ca(C2H3O2)2  that exists   

        iii.  the number of Ca(C2H3O2)2 molecules present


1
Expert's answer
2021-01-20T03:17:16-0500

i.) Molecular mass of"Ca(C_2H_3O_2)_2"

"=40+(12.0102*2+1.00794*3+15.9994*2)2\newline =158.166g/mol"


ii.) No. of moles"=\dfrac{mass}{molar mass}"



"Moles=\dfrac{200}{158.166}"


"=1.2645mol"

iii.) Number of molecules present.

Avogadro's number="6.0221*10^{23}" molecules per mole.

"1mol=6.0221*10^{23}\newline 1.2645=??"

"=1.2645×6.0221*10^{ 23}\newline =7.6149*10^{23} molecules"